SpaceRef

SpaceRef


ISS On-Orbit Status 14 Apr 2002

Status Report From: NASA HQ
Posted: Sunday, April 14, 2002

All ISS systems continue to function nominally, except as noted previously or below.  Flight Day 7 of Mission 8A.

The crews were congratulated on yesterday's "awesome" EVA-2 which set the stage for today's activities.  Major highlight today: the successful completion of the third spacewalk, this time again by MS5 Steve Smith (EV1) and MS1 Rex Walheim (EV2).

EVA-3 Summary:  EVA-3 began at 9:48 am EDT, 45 min. ahead of schedule.  Since it involved rewiring of SSRMS/Canadarm2 connections, CDR Bloomfield and PLT Frick used the Shuttle RMS, fitted with a work platform, for "flying" EV1.  Activities proceeded smoothly.  First task was the release of the LCA (Lab cradle assembly) claw, now no longer needed to hold the S-Zero truss to the Lab.  Most of the time was spent on new installation of J300 panel connectors and subsequent reconfiguration of J400 PDGF (power and data grapple fixture) connectors, all of which provides a new route for power, computer data and video flow to the SSRMS/Canadarm2 when it is attached on the MT (mobile transporter) PDGF after Mission STS-111 (UF-2) in June.  For these tasks, MCC-H commanded powerdowns, and Dan Bursch performed internal Lab J400 panel reconfiguration (for which -- as had been determined earlier -- the TeSS [temporary sleep station] did not have to be removed).  Smith and Walheim then spent about an hour on the MT (12:30 pm), removing launch restraints on the UMA (umbilical mechanism assembly), releasing launch locks on the LDU (linear drive unit) and RSU (roller suspension unit) and removing a small thermal cover on the MT's RPCM (remote power control module).  They also completed the installation of the two CIDs (circuit interrupt devices) #7 and #8, deferred from EVA-1 on 4/11.  They then finished work on the J400 power reconfiguration, so that the SSRMS and S0 could be brought back to their standard powered status later tonight by MCC-H (early tests already indicate good video connections).  Two remaining tasks were deferred to EVA-4 (4/16): transfer of tools and circuit testing devices from the Atlantis to a stowage bag outside the Node (they were temporarily wire-tied to the exterior), and the installation of an EVA translation aid called Airlock Spur, a 4 m long beam fitted with handrails to provide a short-cut for spacewalkers between the Airlock (A/L) and the forward side of the S0 truss and the Lab.  The EVA ended at 4:15 pm, after a total duration of 6 h 27 min.  It was the 37th ISS spacewalk, bringing total assembly EVA time to 229 h 50 min, and the 12th EVA conducted from the ISS itself (total time: 62 h 35 min).  Steve Smith has now performed seven spacewalks, one less than the current record holder, Jerry Ross (51 h 41 min).

During yesterday's EVA-2 a safing bolt of the TUS 2 IUA (trailing umbilical system #2/interface umbilical assembly) did not release. With this bolt in, the TUS 2 cable cutter cannot be actuated, but the loss of this capability is not affecting current MT operations, and troubleshooting was therefore not addressed during today's EVA.  It is expected that the issue will be investigated/repaired during EVA-4. [The cable cutter assembly in each of the two IUAs allows for remote disconnection of a TUS cable in a contingency situation.  If a reel assembly fails or a TUS cable jams, the TUS cable can be severed to allow the MT railcar to safely reach the nearest worksite and receive power to prevent freezing hardware.]

Meanwhile on board the station, CDR Yuri Onufrienko worked in the FGB, replacing dust collector fans as part of regular maintenance. He also completed daily routine systems maintenance on the SOSH life support system, prepared the daily "delta" file for the IMS (inventory management system) and took counter readings of the SVO water supply and SP toilet flush systems for read-down to MCC-M.

FE-2 Dan Bursch completed the ADVASC (advanced astroculture) cable disconnections and transfer procedure, which configured the EXPRESS Rack 4 (ER4) for the subsequent CGBA (commercial generic bioprocessing apparatus) transfer from the Shuttle middeck and installation in ER4 Locker 3, to be completed by FE-1 Carl Walz late in the day.  Bursch stowed the ADVASC units temporarily until they can be transferred to the Shuttle.

Bursch also completed another gas calibration and sample collection on the BPS (biomass processing system) test facility in three parts, today on plant growth chamber #1.

The EVARM (EVA radiation monitoring) activity continued, with EVARM badges carried by the EVA-3 crewmembers in their LCVG (liquid cooling ventilation garment) pockets during the spacewalk.

Transfer of the Lab BTR (biotechnology refrigerator) contents to the Shuttle middeck BTR was to be performed at about 5:00 pm tonight. The BTR was activated by MS4 Ross at 2:20 pm which gave it sufficient time to cool down before Bursch transfers the BTR material from the ISS.

With all four S0 MTS struts firmly installed, all constraints on crew exercising have been lifted, and physical exercise is again being scheduled for all crewmembers.

At 5:59 am EDT, attitude control was handed over to Atlantis, for maneuvering, at 6:39 am, to optimum attitude for dumping waste water overboard.  The contents of the Orbiter's waste water tank, two PWRs (portable water reservoirs) from the ISS and a CWC (collapsible water container) with Shuttle condensate were emptied into space from the Orbiter vent nozzle.  The venting was planned such that it protected an additional docked day without requiring additional water dumps. Attitude control returned to the station CMGs at 8:45 am.

Later, at 4:54 pm, the Orbiter again assumed control over the stack, to steer it to the proper attitude for the second reboost maneuver, scheduled to begin at 5:44 pm for one hour, with computer-controlled and structurally "detuned" jet firings. Expected altitude increase: 1 s.mi. Yesterday's reboost #1 was successfully performed at the scheduled time (6:39 pm), resulting in a mean altitude increase of 1.6 km (1.0 s.mi.). [Orbital data after Reboost-1 see below.]

Today's target areas of the CEO (crew earth observation) and Shuttle earth imaging programs were Mediterranean Dust (crew to look to the east of track for dust blowing from Libya over the Mediterranean. Oblique shots to define boundaries), Brazilian Agriculture (weather should have been remarkably clear over this part of Brazil. Pass began a little past Sao Paulo, which was probably be covered in cloud), Saharan Dust (dust from the Sahara was moved by a cold front south over the Sahel and then out over the Atlantic. Crew to look East of track for the distribution and boundaries, with the coast of Guinea providing geographic context; oblique views), Amazon Mouth (this area is a special focus of scientists using astronaut photography to look at sediments leaving the river. Crew was also to photograph island structure in the deltaic area. With luck, daily thunderstorms that form over land were late enough to provide gaps in cloud cover), Caracas to Montserrat (direct pass over Caracas, Venezuela, with likely clear views. Two minutes later ISS passed just N of Montserrat. Soufriere Hills Volcano increased in activity the week of April 6 with pyroclastic flows reaching the sea, so crew was to look for possible steam. The exact location of the Soufriere Hills Volcano is at 16.72 N 62.18 W), London (nadir pass over London and the Thames. Although weather is quite changeable, may have been a low-cloud opportunity for London and surrounding areas), Ohio (this was a good opportunity for Akron [Hudson], OH, a little south of track. Pass began at the junction of the Mississippi and Ohio Rivers. There should have been good opportunities for Cincinnati, Akron, Buffalo, and Montreal. Crew was to continue photographing the Gulf of St. Lawrence and Newfoundland), San Diego to Colorado (this pass began over San Diego, then followed N of the route of the Colorado River. Ski resorts in Aspen and Crested Butte were S of the Colorado River, S of track), Fiji/Kiribati (this pass began over Fiji, which were partly cloud covered. It then crossed the Wallis islands, Tokelau, and Phoenix Islands [Kiribati]. Coral reef mapping), Hawaii (the Big Island will be just N of track. Crew to look at Kilauea activity. An oblique view could highlight any volcanic aerosols), Great Barrier Reef (part of the Southern Great Barrier Reef features large clusters of small reefs in the Swain and Capricorn Groups. As ISS left the coast, crew was to photograph the bays and their sediment loads N of track. Mapping images of the reefs).

ISS Orbit (as of this last night, 00:31 am EDT  after reboost 1):
Mean altitude -- 386.7 km
Apogee -- 388.9 km
Perigee -- 384.6 km
Period -- 92.3 min.
Inclination (to Equator) -- 51.64 deg
Eccentricity -- 0.0003172
Orbits per 24-hr. day -- 15.60
Altitude increase -- 1600 m (mean) in last 24 hours
Revolutions since FGB/Zarya launch (Nov. =9298) -- 19404
Current Flight Attitude -- LVLH (local vertical/local horizontal =3D 'earth-fixed': z-axis in local vertical, x-axis in velocity vector [yaw: -10 deg, pitch: -7 deg., roll: 0 deg]).

// end //

More status reports and news releases or top stories.

Please follow SpaceRef on Twitter and Like us on Facebook.

SpaceRef Newsletter